thatguyaj44:

A point Q on a segment with endpoints A (2, −1) and C (4, 2) partitions the segment in a 4:1 ratio. Find Q. You must show all work to receive credit.

jhonyy9:

like a first idea use the distance formula @Laylalyssa

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 like a first idea use the distance formula @supie \(\color{#0cbb34}{\text{End of Quote}}\)

supie:

The distance formula is \(d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\) ??

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie The distance formula is \(d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\) ?? \(\color{#0cbb34}{\text{End of Quote}}\) yes supie using this you can find the segment what you need partitions it in the ratio of 4:1

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie The distance formula is \(d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\) ?? \(\color{#0cbb34}{\text{End of Quote}}\) good job

supie:

Thanx So then we plug in x2, x1, y2, and y1 x2=4 x1=2 y2=2 y1=-1 so \(d=\sqrt{(4-2)^2+(2-(-1))^2}\) ???

YECR:

do u still ned help?